5.3 Quality control. As part of a quality control process for computer chips, an engineer at a factory randomly samples 212 chips during a week of production to test the current rate of chips with severe defects. She finds that 27 of the chips are defective.

  1. (a) What population is under consideration in the data set?

  2. (b) What parameter is being estimated?

  3. (c) What is the point estimate for the parameter?

  4. (d) What is the name of the statistic can we use to measure the uncertainty of the point estimate?

  5. (e) Compute the value from part (d) for this context.

  6. (f) The historical rate of defects is 10%. Should the engineer be surprised by the observed rate of defects

    during the current week?

  7. (g) Suppose the true population value was found to be 10%. If we use this proportion to recompute the

    value in part (e) using p = 0.1 instead of pˆ, does the resulting value change much?
Answer #1

a)

population is chips produced

b)

parameter is proportion of defective chips

c)

p^ = X/n = 27/212 = 0.1274

d)

standard error

e)

standard error of proportion =sqrt(pq/n) = sqrt(0.1274* (1-0.1274)/212) = 0.0229

Please rate

Solved first four sub-parts by policy, post rest parts again